# Learning to Refuse: Towards Mitigating Privacy Risks in LLMs

### v1

@@ 1 Introduction
Figure 1: The example of extracting private information from LLMs.
@@
Considering these problems, we propose [Uncaptioned image] RETURN, a Real-world pErsonal daTa UnleaRNing dataset. As illustrated in Figure 2, we collect extensive background information on celebrities from Wikipedia and use GPT-4 Achiam et al. (2023) to generate 20$\times$×QA pairs for each individual. After manual and automated validation, we obtain a dataset of 2,492 individuals, each with a (Name, 20$\times$×QA pairs) data instance. Next, we could select a base model to evaluate the MU methods on [Uncaptioned image] RETURN. In this work, we take LLaMA-3-8B-Instruct AI@Meta (2024) as an example. We first identify individuals with deep memorization in the model and then divide them into the forget set and the retain set. Our goal is for the model to protect the information of individuals in the forget set, ensuring that questions related to these individuals are not answered correctly, while maintaining the model’s performance on the retain set.
Existing MU methods often suffer from sensitivity to hyperparameter selection or the inability to effectively distinguish between the forget set and the retain set. To mitigate the drawbacks of existing methods, we propose a simple yet novel unlearning method: Name-Aware Unlearning Framework (NAUF) for privacy protection. The framework comprises two key components: Name-Aware Refusal Answer and Contrastive Data Augmentation. The Name-Aware Refusal Answer is designed to help the model learn which individuals’ information should be protected, and the Contrastive Data Augmentation aims to expand the distribution of both the forget set and the retain set for enhancing the generalization of our method. We evaluate the effectiveness of our proposed method on [Uncaptioned image] RETURN and compare it with the baseline methods, and the results show that our proposed NAUF achieves a state-of-the-art average unlearning score, outperforming the best baseline method by 5.65 points.
@@
2 Related Work
Memorization and Privacy Risks of LLMs.
Machine Unlearning for LLMs.
@@ 3.1 Data Construction
We begin by leveraging PopQA Mallen et al. (2022) to collect a large set of names of individuals. PopQA is a large-scale open-domain question-answering (QA) dataset constructed by Mallen et al. (2022), consisting of 14k entity-centric QA pairs. Each pair comes with the original [subject entity, relationship type, object entity] annotation, as well as Wikipedia monthly page views. Specifically, for the data in PopQA, we collect “subject entity” if the “relationship type” is within [ occupation, place of birth, father, mother]; and we collect “object entity” if the “relationship type” is within [ producer, director, screenwriter, composer, author].
After gathering these names, we retrieve their corresponding Wikipedia pages and extract the abstracts from these pages as background information [^2] https://github.com/martin-majlis/Wikipedia-API. We then filter the background information to retain only those whose word count falls between 100 and 500 words. Through this process, we ultimately obtain 2,516 records consisting of (Name, Background Information). Next, given each pair of name and the background information, we use a prompt to generate 20$\times$×QA pairs with GPT4 Achiam et al. (2023). The prompt template is shown in Table 3 in the Appendix.
@@ 3.2 Identifying Individuals with Deep Memorization
To perform unlearning on LLaMA-3, we first need to identify which individuals the model has deeply memorized. We ask the model to answer the questions for each individual in [Uncaptioned image] RETURN, then calculate the average accuracy by comparing the model’s predicted answers with the gold answers using a Natural Language Inference (NLI) model  [^3] We choose deberta-v3-base-tasksource-nli Sileo (2023) to evaluate the correctness of model’s prediction.. If the prediction is "entailment" or "neutral," we consider the model’s answer correct; if the NLI model’s prediction is "contradiction," we consider the model’s answer incorrect [^4] When the model’s predicted answer is partially correct and does not conflict with the gold answer, the NLI model’s prediction is ”neutral.” Therefore, we will consider ”neutral” as correct in this context.. The accuracy distribution of LLaMA-3 on [Uncaptioned image] RETURN is shown in Figure 3. The higher the accuracy, the more deeply the model memorizes the individual’s information. Finally, we take 466 individuals with accuracy $\geq$ 0.8 as individuals with deep memorization for the subsequent unlearning experiments.
@@ 3.4 Evaluation Metrics
4 Machine Unlearning for LLMs
In this section, we introduce MU baselines. Then we propose a novel method for mitigating privacy risks in LLMs: Name-Aware Unlearning Framework for Privacy Protection.
@@ 4.1 Baseline Methods
A typical MU method often have two parts: unlearning on the forget set and regularization on the retain set. These two types of loss can be used in any combination.
4.1.1 Unlearning on Forget Set
Gradient Ascent.
Negative Preference Optimization.
$\mathcal{L}_{NPO}(\mathcal{D}^{F},\mathcal{M}_{u},\mathcal{M}_{o})$ (4)
Relabeled Gradient Descent.
$\mathcal{L}_{RGD}(\mathcal{D}^{F},\mathcal{M}_{u})$ (5)
Relabeled Direct Preference Optimization.
$\mathcal{L}_{RDPO}(\mathcal{D}^{F},\mathcal{M}_{u},\mathcal{M}_{o})$ (6)
$\hfill-\beta\log\frac{\mathcal{M}_{u}(y|x)}{\mathcal{M}_{o}(y|x)})]$
4.1.2 Regularization on Retain Set
Gradient Descent Regularization.
$\mathcal{L}_{GD}(\mathcal{D}^{R},\mathcal{M}_{u})$ (7)
Kullback-Leibler Divergence Regularization.
$\mathcal{L}_{KL}(\mathcal{D}^{R},\mathcal{M}_{u},\mathcal{M}_{o})$ (8)
4.2 Name-Aware Unlearning Framework
In our pilot study, we find that RGD achieves the comparatively best performance in protecting the privacy of individuals in the forget set. However, the model’s performance on the retain set is significantly affected. The model tends to refuse to answer the questions on the retain set, which is not desirable. To address this issue, we propose a novel method: Name-Aware Unlearning Framework (NAUF) for privacy protection. The framework comprises two key components: Name-Aware Refusal Answer and Contrastive Data Augmentation.
Name-Aware Refusal Answer.
First, we relabel the questions in the forget set with a name-aware refusal answer, such as "I’m afraid I can’t help with inquiries about NAME." Then we could perform gradient ascent on the loss over the relabeled forget set. The name-aware refusal answer is designed to help the model learn which individuals’ information should be protected. We curate 100 name-aware refusal answer templates $\mathcal{D}^{refuse}$ using GPT-4, and some examples are shown in Table 4 in the Appendix.
Contrastive Data Augmentation.
In addition, given the limited number of QA pairs for each individual, we propose contrastive data augmentation (CDA) to enhance the generalization of unlearning. Specifically:
  • For each individual in the forget set, we randomly sample questions from other individuals in the forget set and replace the name with the target individual’s name. Then relabel the questions with the name-aware refusal answer.
  • For each individual in the retain set, we also randomly sample questions from other individuals in the forget set and replace the name with the target individual’s name, but we use the original model’s prediction for that question as the relabeled answer.
This contrastive data augmentation strategy expands the distribution of both the forget set and the retain set. For simplicity, we expand the forget set and the retain set by doubling the amount of data.
5 Experiments
@@ 5.1 Implementation Details
The $\beta$ for NPO and RDPO is set to 0.1. We use the AdamW optimizer with a learning rate of 1e-5 for all experiments. We set the batch size to 32 and train the model for 5 epochs. Considering the computational budget, we constraint that the number of samples used from the retain set is equal to the number of the entire forget set in each epoch. All experiments are conducted with 2 NVIDIA A100 (40GB) GPUs.
@@ 5.2 Main Results
Table 2: The main results of the experiments. Forget S. denotes Forget Score, Retain S. denotes Retain Score, WG denotes WinoGrande, LQA denotes LogiQA, LAM denotes LAMBADA. Oracle refers to using the original model directly to compute the metrics without applying any unlearning. Notably, NS denotes "NonSense", which means the model’s prediction is meaningless, and we take it as 0 for computing the average. We highlight the best results in bold.
@@ 5.3 Analysis
6 Conclusion
@@
The proposed NAUF method is designed for individual-level privacy protection. Our goal is to protect all information about an individual, ensuring that the model refuses to answer any questions related to that individual. However, this method does not provide fine-grained protection of the target individual’s information. In other words, it cannot distinguish between questions that can be answered and those that are too sensitive to answer. Future work will explore how to align the model with human judgment, enabling it to discern which personal information can be publicly discussed and which information, potentially susceptible to malicious use, should be protected.
Appendix A QA Pairs Generation Template
@@
Appendix B Examples of Name-Aware Refusal Answer Templates
We provide 10 examples of name-aware refusal answer templates in Table 4. It is designed to help LLM learn which individuals’ information should be protected.
@@ Appendix B Examples of Name-Aware Refusal Answer Templates
Name-Aware Refusal Answers Apologies, I cannot discuss [NAME]-related inquiries. Sorry, I’m unable to provide information regarding [NAME]. Unfortunately, I’m not able to address questions about [NAME]. I’m sorry, but I can’t offer insights on [NAME]. Regrettably, [NAME] is not a topic I can discuss. I’m unable to engage in discussions about [NAME]. Sorry, I’m not equipped to answer queries about [NAME]. I can’t provide information on topics involving [NAME]. Unfortunately, [NAME] falls outside the scope of my responses. Apologies, but I’m unable to delve into matters concerning [NAME].
@@
Table 4: Name-aware refusal answer templates.

### v2

@@ 1 Introduction
Figure 1: The example of extracting private information from LLMs. When an individual practices RTBF, the model should protect his/her private information.
@@
Considering these problems, we propose [Uncaptioned image] RETURN, a Real-world pErsonal daTa UnleaRNing dataset. As illustrated in Figure 2, we collect extensive background information on celebrities from Wikipedia and use GPT-4 Achiam et al. (2023) to generate 20$\times$×QA pairs for each individual. After manual and automated validation, we obtain a dataset of 2,492 individuals, each with a (Name, 20$\times$×QA pairs) data instance. Next, we could select a base model to evaluate the MU methods on this dataset. In this work, we take LLaMA-3-8B-Instruct AI@Meta (2024) as an example. We first identify individuals with deep memorization in the model and then divide them into the forget set and the retain set. Our goal is for the model to protect the information of individuals in the forget set, ensuring that questions related to these individuals are not answered correctly, while maintaining the model’s performance on the retain set.
Existing MU methods often face challenges. One category, based on gradient ascentLiu et al. (2024a), suffers from sensitivity to hyperparameter selection or inability to effectively distinguish between the forget set and the retain set. Another category transforms traditional gradient ascent into gradient descent on a relabeled forget set, such as Relabeled Gradient Descent (RGD) Maini et al. (2024). By training the model to generate uninformed answers like "I don’t know", RGD achieves better performance in protecting the privacy of individuals in the forget set. However, our pilot study finds that RGD significantly affects the model’s performance on the retain set, causing the model to refuse to answer questions it should address. To overcome these limitations, we propose a simple yet novel unlearning method: Name-Aware Unlearning Framework (NAUF) for privacy protection. The framework comprises two key components: Name-Aware Refusal Answer and Contrastive Data Augmentation. The Name-Aware Refusal Answer is designed to help the model learn which individuals’ information should be protected, and the Contrastive Data Augmentation aims to expand the distribution of both the forget set and the retain set for enhancing the generalization of our method. We evaluate the effectiveness of NAUF on our proposed dataset and compare it with the baseline methods, and the results show that our proposed NAUF achieves a state-of-the-art average unlearning score, outperforming the best baseline method by 5.65 points.
@@ 2.1 Data Construction
We begin by leveraging PopQA Mallen et al. (2022) to collect a large set of names of individuals. PopQA is a large-scale open-domain question-answering (QA) dataset constructed by Mallen et al. (2022), consisting of 14k entity-centric QA pairs. Each pair comes with the original [subject entity, relationship type, object entity] annotation, as well as the Wikipedia monthly page views for both the subject and object entities, which serve as a measure of their popularity. Specifically, for the data in PopQA, we collect “subject entity” if the “relationship type” is within [occupation, place of birth, father, mother]; and we collect “object entity” if the “relationship type” is within [producer, director, screenwriter, composer, author].
After gathering these names, we retrieve their corresponding Wikipedia pages and extract the abstracts from these pages as background information [^2] https://github.com/martin-majlis/Wikipedia-API. We then filter the background information to retain only those whose word count falls between 100 and 500 words. Through this process, we ultimately obtain 2,516 records consisting of (Name, Background Information). Next, given each pair of name and the background information, we use a prompt to generate 20$\times$×QA pairs with GPT4 Achiam et al. (2023). The prompt template is shown in Appendix D.
@@ 2.2 Identifying Individuals with Deep Memorization
To perform unlearning on LLaMA-3, we first need to identify which individuals the model has deeply memorized. We ask the model to answer the questions for each individual in the dataset, then calculate the average accuracy by comparing the model’s predicted answers with the gold answers using a Natural Language Inference (NLI) model  [^3] We choose deberta-v3-base-tasksource-nli Sileo (2023) to evaluate the correctness of model’s prediction.. If the prediction is "entailment" or "neutral," we consider the model’s answer correct; if the NLI model’s prediction is "contradiction," we consider the model’s answer incorrect [^4] When the model’s predicted answer is partially correct and does not conflict with the gold answer, the NLI model’s prediction is ”neutral.” Therefore, we will consider ”neutral” as correct in this context.. The accuracy distribution of LLaMA-3 on [Uncaptioned image] RETURN is shown in Figure 3. The higher the accuracy, the more deeply the model memorizes the individual’s information. Finally, we take 466 individuals with accuracy $\geq$ 0.8 as individuals with deep memorization for the subsequent unlearning experiments.
We analyze the popularity of individuals in our dataset, categorized by those with and without LLaMA-3’s deep memorization. We find that there is a significant difference in average popularity: 68620.9 for individuals with deep memorization versus 36841.1 for those without. This may be because highly popular individuals tend to have more diverse and detailed information available online, such as biographical details, interviews, news coverage, and social media activity, thus increasing the likelihood of deep memorization.
@@ 2.4 Evaluation Metrics
3 Name-Aware Unlearning Framework
Existing MU methods often face challenges in effectively protecting privacy in the forget set while maintaining model performance on the retain set. To address these challenges, we propose a novel method: Name-Aware Unlearning Framework (NAUF) for privacy protection. The framework comprises two key components: Name-Aware Refusal Answer and Contrastive Data Augmentation.
Name-Aware Refusal Answer.
First, we relabel the questions in the forget set with a name-aware refusal answer, such as "I’m afraid I can’t help with inquiries about NAME." Then we could perform gradient descent on the loss over the relabeled forget set. The name-aware refusal answer is designed to help the model learn which individuals’ information should be protected. We curate 100 name-aware refusal answer templates $\mathcal{D}^{refuse}$ using GPT-4, which are shown in Appendix E.
Figure 4: The example of CDA for an individual in the forget set. Here we take Darrell Hammond as target individual.
Contrastive Data Augmentation.
In addition, due to the limited number of QA pairs available for each individual, we propose contrastive data augmentation (CDA) as a straightforward and cost-effective method to enhance the quantity and diversity of data. This approach aims to improve the model’s ability to generalize across information related to the targeted individuals. Specifically:
  • For each individual in the forget set, we randomly sample questions from other individuals in the forget or retain set and replace the name with the target individual’s name. Then relabel the questions with the name-aware refusal answer. An example is shown in Figure 4.
  • For each individual in the retain set, we also randomly sample questions from other individuals in the forget or retain set and replace the name with the target individual’s name. Then we input the modified questions into the original model, and use the original model’s prediction for that question as the relabeled answer. An example is shown in Figure 5.
This contrastive data augmentation strategy expands the distribution of both the forget set and the retain set, and subsequent experiments demonstrate that it significantly improves the performance of our proposed method. For simplicity, we expand the forget set and the retain set by doubling the amount of data.
Figure 5: The example of CDA for an individual in the retain set. Here we take Brian Eno as target individual.
4 Experiments
@@ 4.1 Baseline Methods
A typical MU method generally consists of two components: unlearning on the forget set and regularization on the retain set. These two types of loss can be used in any combination.
Unlearning on Forget Set:
The unlearning process on the forget set includes methods such as Gradient Ascent (GA), Negative Preference Optimization (NPD), Relabeled Gradient Descent (RGD), and Relabeled Direct Preference Optimization (RDPO). The details of these methods are available in Appendix B.
Regularization on Retain Set.
The regularization methods on the retain set include Gradient Descent (GD) regularization and Kullback-Leibler Divergence (KLR) regularization . The details of these regularization methods are available in Appendix C.
@@ 4.2 Implementation Details
The $\beta$ for NPO and RDPO is set to 0.1. We use the AdamW optimizer with a learning rate of 1e-5 for all experiments. We set the batch size to 32 and train the model for 5 epochs. Considering the computational budget, we constraint that the number of samples used from the retain set is equal to the number of the entire forget set in each epoch. All experiments are conducted with 2 NVIDIA A100-40GB GPUs, and each take approximately 1-2 hours with Deepspeed Zero3 Offload.
@@ 4.3 Main Results
Table 2: The main results of the experiments. Forget S. denotes Forget Score, Retain S. denotes Retain Score, WG denotes WinoGrande, LQA denotes LogiQA, LAM denotes LAMBADA. Oracle refers to using the original model directly to compute the metrics without applying any unlearning. Notably, NS denotes "NonSense", which means the model’s prediction is meaningless, and we take it as 0 for computing the average. We highlight the best results in bold, the second highest in underline.
@@ 4.4 Analysis
Method Forget Score Retain Score Avg.
Forget:Retain = 1:99
RGD 96.67 22.89 59.78
NAUF 93.33 28.46 60.90
Forget:Retain = 5:95
RGD 86.87 26.19 56.53
NAUF 94.95 67.59 81.27
Forget:Retain = 10:90
RGD 96.21 52.01 74.11
NAUF 93.69 67.82 80.76
Forget:Retain = 20:80
RGD 43.02 75.08 59.05
NAUF 93.53 71.32 82.43
Table 3: Unlearning Performance of RGD/NAUF with KLD Regularization across Different Ratio between Forget Set and Retain Set.
Unlearning Performance across Different Ratio between Forget Set and Retain Set.
We conduct additional analyze the impact of different data ratios on the MU algorithms. As shown in Table 3, the results demonstrate that increasing the proportion of the forget set could improve the retain score, which because we constraint the number of samples used from the retain set is equal to the number of the entire forget set in each epoch. We also find that our method consistently maintains high forget scores (90+).
Figure 8: Average unlearning score of NAUF with KLD Regularization across different numbers of augmented data.
Average Unlearning Score of Different CDA Number.
We conducted experiments to assess the impact of different number of augmented data on the average unlearning score, as shown in Figure 8. The results indicate that as the number of augmented data increases, the performance gradually improves, reaching its peak when the augmented data count reaches 40. This suggests that appropriate data augmentation can enhance unlearning performance.
5 Conclusion and Future Work
@@
This study focuses on individual-level privacy protection through a name-aware unlearning framework. To broaden this approach to other types of sensitive data, future work could generalize the protection to the entity level or concept level. Such a modification would enable the model to learn to refuse instructions related to specific entities—like anime characters—or concepts such as locations in copyrighted books. These adaptations would enhance the framework’s versatility and applicability to a wider range of privacy concerns.
@@
The Size of Dataset.
The proposed [Uncaptioned image] RETURNdataset is constructed based on PopQA, containing a total of 2,492 entries. Technically, extracting data directly from Wikipedia to construct a larger dataset is feasible. However, due to our limited resources, we cannot afford the costs associated with GPT-4 api for constructing QA pairs. Therefore, we left the development of a larger scale dataset as future work.
Fine-grained Protection.
The current work is focused on exploring whether a model can protect all information about an individual based on partial data, thereby maximizing privacy security for that individual. However, this method does not provide fine-grained protection of the target individual’s information. Future work could explore fine-grained protection of the target individual’s information. The goal is to enable the model to autonomously discern which pieces of information might be exploited for harmful purposes and therefore should be protected, without compromising the accessibility of benign information.
Appendix A Related Work
Memorization and Privacy Risks of LLMs.
Machine Unlearning for LLMs.
Appendix B Unlearning on Forget Set
Gradient Ascent.
Negative Preference Optimization.
$\mathcal{L}_{NPO}(\mathcal{D}^{F},\mathcal{M}_{u},\mathcal{M}_{o})$ (4)
Relabeled Gradient Descent.
$\mathcal{L}_{RGD}(\mathcal{D}^{F},\mathcal{M}_{u})$ (5)
Relabeled Direct Preference Optimization.
$\mathcal{L}_{RDPO}(\mathcal{D}^{F},\mathcal{M}_{u},\mathcal{M}_{o})$ (6)
$\hfill-\beta\log\frac{\mathcal{M}_{u}(y|x)}{\mathcal{M}_{o}(y|x)})]$
Appendix C Regularization on Retain Set
Gradient Descent Regularization.
$\mathcal{L}_{GD}(\mathcal{D}^{R},\mathcal{M}_{u})$ (7)
Kullback-Leibler Divergence Regularization.
$\mathcal{L}_{KL}(\mathcal{D}^{R},\mathcal{M}_{u},\mathcal{M}_{o})$ (8)
Appendix D QA Pairs Generation Template
@@
Appendix E Examples of Name-Aware Refusal Answer Templates
@@
We provide the name-aware refusal answer templates in Table 5, Table 6 and Table 7. It is designed to help LLM learn which individuals’ information should be protected.
Name-Aware Refusal Answers Apologies, I cannot discuss [NAME]-related inquiries. Sorry, I’m unable to provide information regarding [NAME]. Unfortunately, I’m not able to address questions about [NAME]. I’m sorry, but I can’t offer insights on [NAME]. Regrettably, [NAME] is not a topic I can discuss. I’m unable to engage in discussions about [NAME]. Sorry, I’m not equipped to answer queries about [NAME]. I can’t provide information on topics involving [NAME]. Unfortunately, [NAME] falls outside the scope of my responses. Apologies, but I’m unable to delve into matters concerning [NAME]. Sorry, I can’t assist with questions related to [NAME]. I’m afraid I can’t help with inquiries about [NAME]. Unfortunately, [NAME] is off-limits for discussion. Apologies, [NAME]-related questions are not within my capabilities. I’m unable to provide any information about [NAME]. Sorry, discussions about [NAME] are not possible. Regrettably, [NAME] is not a subject I can address. I’m sorry, but I’m not able to discuss [NAME]. Unfortunately, I can’t engage in conversations involving [NAME]. Sorry, but I cannot discuss topics pertaining to [NAME].
Table 5: Name-aware refusal answer templates (1-20).
Name-Aware Refusal Answers Apologies, [NAME]-related queries are beyond my purview. Sorry, but I’m unable to provide insights on [NAME]. Regrettably, [NAME] is a topic I’m not allowed to discuss. I’m sorry, but I cannot provide any information about [NAME]. Unfortunately, [NAME] is not a subject I can provide details on. Apologies, but I’m not authorized to discuss matters regarding [NAME]. Sorry, but [NAME] is outside the realm of my knowledge. Regrettably, I cannot offer any insights into [NAME]. I’m sorry, but I’m not permitted to discuss [NAME]. Unfortunately, [NAME] is not a topic I’m able to address. Apologies, but discussions about [NAME] are not permitted. Sorry, but I’m not allowed to provide information about [NAME]. Regrettably, I’m not authorized to discuss [NAME]. I’m sorry, but [NAME] is not within the bounds of this conversation. Unfortunately, [NAME] is a topic I’m not permitted to discuss. Apologies, but [NAME]-related matters are not within my jurisdiction. Sorry, but I’m not able to provide any insights on [NAME]. Regrettably, I cannot provide information about [NAME]. I’m sorry, but I’m unable to discuss [NAME]. Unfortunately, [NAME] is not a topic I can elaborate on. Apologies, but I cannot engage in discussions about [NAME]. Sorry, but I’m not authorized to address questions about [NAME]. Regrettably, [NAME] is a topic I’m not allowed to discuss. I’m sorry, but I cannot offer any insights into [NAME]. Unfortunately, [NAME] is off-limits for discussion. Apologies, but [NAME] is not a subject I can discuss. Sorry, but I’m not permitted to provide information about [NAME]. Regrettably, I’m unable to address queries about [NAME]. I’m sorry, but discussions about [NAME] are not allowed. Unfortunately, [NAME] is beyond the scope of this conversation. Apologies, but I cannot discuss matters concerning [NAME]. Sorry, but [NAME] is not a topic I’m allowed to discuss. Regrettably, [NAME] is not something I can provide information on. I’m sorry, but I cannot engage in discussions regarding [NAME]. Unfortunately, [NAME] is outside the bounds of my knowledge. Apologies, but I’m not authorized to provide insights on [NAME]. Sorry, but [NAME]-related inquiries cannot be addressed. Regrettably, [NAME] is a topic I’m not permitted to discuss. I’m sorry, but I’m not authorized to discuss matters pertaining to [NAME]. Unfortunately, I cannot provide any information about [NAME].
Table 6: Name-aware refusal answer templates (21-60).
Name-Aware Refusal Answers Apologies, but discussions involving [NAME] are not allowed. Sorry, but I cannot discuss [NAME]-related topics. Regrettably, I’m not permitted to address questions about [NAME]. I’m sorry, but I’m unable to offer insights into [NAME]. Unfortunately, [NAME] is a topic that I cannot discuss. Apologies, but [NAME] is not within my area of expertise. Sorry, but I cannot provide any information on [NAME]. Regrettably, discussions about [NAME] are not permitted. I’m sorry, but I’m not authorized to discuss [NAME]. Unfortunately, [NAME] is not something I can address. Apologies, but I cannot engage in discussions about [NAME]. Sorry, but I’m not allowed to provide insights on [NAME]. Regrettably, [NAME] is off-limits for discussion. I’m sorry, but I’m unable to discuss matters pertaining to [NAME]. Unfortunately, I cannot provide any information regarding [NAME]. Apologies, but discussions about [NAME] are not permitted. Sorry, but I’m not authorized to discuss [NAME]. Regrettably, [NAME] is a topic I’m not allowed to address. I’m sorry, but I cannot provide insights into [NAME]. Unfortunately, [NAME] is outside the scope of this conversation. Apologies, but I cannot discuss matters related to [NAME]. Sorry, but I’m not permitted to provide information about [NAME]. Regrettably, I’m unable to address queries about [NAME]. I’m sorry, but I cannot engage in discussions involving [NAME]. Unfortunately, [NAME] is not a topic I’m authorized to discuss. Apologies, but I’m not allowed to discuss [NAME]-related inquiries. Sorry, but I cannot provide information on topics concerning [NAME]. Regrettably, discussions about [NAME] are not within my jurisdiction. I’m sorry, but I’m not permitted to discuss matters regarding [NAME]. Unfortunately, [NAME] is a topic I’m not able to address. Apologies, but I cannot provide any insights on [NAME]. Sorry, but I’m not authorized to discuss matters concerning [NAME]. Regrettably, [NAME] is off-limits for discussion. I’m sorry, but I cannot discuss [NAME]-related queries. Unfortunately, discussions about [NAME] are not allowed. Apologies, but [NAME] is not a subject I can discuss. Sorry, but I cannot engage in discussions about [NAME]. Regrettably, I’m not allowed to provide information about [NAME]. I’m sorry, but I cannot address questions about [NAME]. Unfortunately, [NAME] is not a topic I’m able to discuss.
Table 7: Name-aware refusal answer templates (61-100).